Please i need to expand these two equations (1). (a - 4) (a - 3) (2). (2x + 3) (2x - 3)
Answer by Cromlix(4381)

(a - 4)(a - 3) = a^2 - 7a + 12
(2x + 3)(2x - 3) = 4x^2 - 9.
Hope this helps.
:-)